OpenStudy (anonymous):

Which equation represents a direct linear variation? A. y = x + 2 B. y = ½x C. y=1/x D. y = x2

OpenStudy (anonymous):

B= 1/2x

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

youre welcome anymore questions

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

look, you can close this thread and open a new one but tag me tho

OpenStudy (anonymous):

Which is the direct linear variation equation for the relationship? y varies directly with x and y = 12 when x = 4. A. y = x – 8 B. y = 2x + 4 C. y = 3x D. y = x + 8

OpenStudy (anonymous):

i think its c not sure tho :/

OpenStudy (anonymous):

k thanks again thats all the questions I have for the moment

OpenStudy (anonymous):

alright :)